# SCINE — Software for Chemical Interaction Networks

### v1

@@ Abstract
The software for chemical interaction networks (SCINE) project aims at pushing the frontier of quantum chemical calculations on molecular structures to a new level. While calculations on individual structures as well as on simple relations between them (e.g., as given by an intrinsic reaction coordinate) have become routine in chemistry, new developments have pushed the frontier in the field to high-throughput calculations. Chemical relations may be created by a search for specific molecular properties in a molecular design attempt or they can be defined by a set of elementary reaction steps that form a chemical reaction network. The software modules of SCINE have been designed to facilitate such studies. The features of the modules are (i) general applicability of the applied methodologies ranging from electronic structure (no restriction to specific elements of the periodic table) to microkinetic modeling (with little restrictions on molecularity), full modularity so that SCINE modules can also be applied as stand-alone programs or be exchanged for external software packages that fulfill a similar purpose (to increase options for computational campaigns and to provide alternatives in case of tasks that are hard or impossible to accomplish with certain programs), (ii) high stability and autonomous operations so that control and steering by an operator is as easy as possible, and (iii) easy embedding into complex heterogeneous environments for molecular structures taken individually or in the context of a reaction network. A graphical user interface unites all modules and ensures interoperability. All components of the software have been made available open source and free of charge.
@@ I Introduction
Within SCINE, we address these challenges by a strictly modular architecture, a loose coupling between the individual modules, and consistent interfaces between them. Individual modules which are only loosely coupled in terms of their source code (i.e., the flow of data between them is clearly defined and limited to only a few instances) are easier to maintain than a complex monolithic architecture. At the same time, this makes developing the software easier and less error-prone, as so-called side effects (i.e., unexpected and unwanted effects upon changing the source code) are largely minimized. Finally, this approach naturally leads to individual modules that are fully functional programs in their own right with useful features. For less complex tasks, it is therefore sufficient to install only a single or a few required modules without having to deal with a particularly complicated software stack.
@@ II Overview of SCINE Modules
Figure 2: SCINE modules and their communication.
@@
The calculations are then carried out by another module called PuffinBensberg et al. (2023c). Puffin retrieves information on new calculations from the database, carries them out, and then writes the results back to the database. For carrying out a calculation, Puffin typically invokes other SCINE modules. For example, it can call SwooseBrunken, Csizi, and Reiher (2021); Brunken and Reiher (2020, 2021) for QM/MM calculations or KiNetXBensberg et al. (2023d); Proppe and Reiher (2019); Bensberg and Reiher (2023a) for microkinetic modeling. For most calculations, however, Puffin relies on ReaDuctBensberg et al. (2023e). The ReaDuct module carries out a wide range of structure-manipulating tasks such as structure minimizations, transition state searches, intrinsic reaction coordinate calculations, B-spline interpolation and optimization of reaction pathsVaucher and Reiher (2018), and Newton trajectory calculationsUnsleber, Grimmel, and Reiher (2022).
@@
Large-scale explorations typically require millions of quantum chemical calculations. Therefore, fast methods with sufficient accuracy are of vital importance. Machine learning potentials can offer both high accuracy and high speed when trained on sufficiently accurate data. However, such potentials require extensive and costly training and often have difficulties learning additional data without “forgetting” previous knowledge. Moreover, most structural descriptors cannot efficiently represent many different atom types. We recently addressed these problems by introducing the new concept of lifelong machine learning potentialsEckhoff and Reiher (2023) implemented in the SCINE module Parrot. Other well-known machine learning potentials such as ANISmith, Isayev, and Roitberg (2017); Devereux et al. (2020) and M3GNetChen and Ong (2022) are also available in Parrot.
@@
Since Chemoton is based on the first principles of quantum mechanics, it is applicable to any chemical system. In practice, this might require specific quantum chemical methods. For example, for very large systems such as proteins, even fast semiempirical methods like the ones provided by Sparrow are typically too time-intensive for large-scale explorations. In this case, one can resort to molecular mechanics calculations. To enable such type of modeling out of the box, the SCINE software stack includes a module called Swoose. Swoose can provide energies and forces from molecular mechanics for any chemical system and run molecular mechanics simulations. If no parameters are available for the system, it can automatically parametrize a system-focused force field from quantum chemical reference calculationsBrunken and Reiher (2020). Large systems for which accurate reference calculations would be too costly are automatically fragmented. Moreover, Swoose is also able to carry out QM/MM calculations, in which a key region within a system is automatically defined and then described by an electronic structure model, while the rest of the system is described by a standard force field (supported by machine learning corrections)Brunken and Reiher (2021).
@@
Open-shell systems often require a multi-configurational wavefunction ansatz to be described qualitatively correctly. Multi-configurational methods such as the density renormalization group (DMRG)White (1992, 1993); Chan et al. (2008); Chan and Zgid (2009); Marti and Reiher (2010); Schollwöck (2011); Chan and Sharma (2011); Wouters and Van Neck (2013); Kurashige (2014); Olivares-Amaya et al. (2015); Szalay et al. (2015); Yanai et al. (2015); Baiardi and Reiher (2020) typically require the definition of a so-called active orbital space on input. However, this active space is often not trivial to determine. Our autoCAS algorithmStein and Reiher (2016); Stein, von Burg, and Reiher (2016); Stein and Reiher (2017), implemented in a module of the same nameStein and Reiher (2019); Bensberg et al. (2023f), provides an automated way to reliably determine the active space of any molecule. Moreover, once the active space is determined, autoCAS can also invoke software to carry out multi-configurational calculations. Besides QCMaquisKeller et al. (2015), which is our software for DMRG calculations and driven by autoCAS via OpenMolcasLi Manni et al. (2023), autoCAS is also interfaced to ChronusQWilliams-Young et al. (2020), to PySCFSun et al. (2020), and SerenityUnsleber et al. (2018). We note that QCMaquis can do much more than just calculate the ground state electronic energy of a molecule. It can also optimize excited statesBaiardi et al. (2019); Baiardi, Kelemen, and Reiher (2022), solve the vibrational Schrödinger equation for anharmonic potential energy surfacesBaiardi et al. (2017); Glaser, Baiardi, and Reiher (2023), and carry out quantum dynamics simulationsBaiardi and Reiher (2019); Baiardi (2021). Furthermore, it features embedding approachesDresselhaus et al. (2015); Hedegård and Reiher (2016), a non-orthogonal state interaction methodKnecht et al. (2016), relativistic calculationsBattaglia, Keller, and Knecht (2018), multi-component DMRGMuolo et al. (2020); Feldmann et al. (2022), and N-electron valence state perturbation theory calculaitonsFreitag et al. (2017).
@@
A central module is HeronMüller et al. (2024); Bensberg et al. (2022), our graphical user interface. It aims to make the entire SCINE software ecosystem easily accessible. Heron provides the ability to browse the contents of a database, visualize and analyze existing reaction networks, and operate Chemoton (for example, with the so-called Steering WheelSteiner and Reiher (2023)). Moreover, it is a graphical user interface for Swoose, ReaDuct, autoCAS, and Sparrow. Together with the latter, it provides the ability to interactively carry out quantum chemical calculations, in which the response of a molecule to structural manipulations can be experienced in real time. This can even be coupled with haptic devices, allowing a person to literally feel the forces acting on atoms in a moleculeMarti and Reiher (2009); Haag and Reiher (2013); Haag et al. (2014); Weymuth and Reiher (2021).
@@ II.1.1 Field of application
MolassemblerBensberg et al. (2023b); Sobez and Reiher (2020) provides characterization, manipulation, and conformer generation for organic and inorganic molecules. In principle, it can handle any organic or inorganic molecule, including haptic and multidentate ligands and almost arbitrary coordination spheres. Its general approach to the characterization of molecules and its broad applicability makes it distinct compared to similar programs such as RDKitRDK or openBabelO’Boyle et al. (2011), which focuses on organic molecules, or molSimplifyIoannidis, Gani, and Kulik (2016); Edholm et al. (2023), AARONGuan et al. (2018) and DENOPTIMFoscato, Venkatraman, and Jensen (2019), which provide utilities to design new molecules, or CRESTPracht, Bohle, and Grimme (2020) for conformer generation.
@@
Molassembler characterizes a molecule’s connectivity by a graph and the local geometry of each atom by polyhedral shapes. In organic chemistry, atoms typically have only up to four bonding partners, leading at most to a tetrahedral local environment. By contrast, transition metal atoms can have significantly more neighbors and diverse coordination geometries. Therefore, Molassembler supports polyhedra for the local geometry up to icosahedra and cuboctahedra and can be applied to the full spectrum of organic and inorganic chemistry. Furthermore, it supports multidentate and haptic ligands. Molassembler differentiates stereoisomers through a ranking algorithm based on generalizing the ranking rules for organic molecules provided by the International Union of Pure and Applied ChemistryFavre and Powell (2013). The abstract molecule representation provided by Molassembler can be manipulated to change the connectivity in the molecule and replace or edit ligands to generate new molecules while retaining any chiral information in the remaining molecule wherever possible. Furthermore, Molassembler can generate conformers by enumerating the rotamers of the molecule, generating initial Cartesian coordinates for the conformer guesses before refining the coordinates.
II.1.2 API
@@ II.4.1 Field of application
SparrowBosia et al. (2023a); Bosia, Weymuth, and Reiher (2022); Bosia et al. (2023b) specializes in ultra-fast quantum chemical calculations with semi-empirical methods, which are essential in applications such as interactive quantum chemistry or large-scale reaction network exploration. Calculations with semi-empirical methods based on the neglect of differential overlap approximation — NDDO methods — as well as with methods based on the density functional tight binding approach — DFTB methods — are available. Ground state calculations yield electronic energies, Hessian matrices, nuclear gradients, bond orders, and thermodynamic properties such as Gibbs free energies. Electronically excited states can also be calculated — in a configuration interaction approach with NDDO methods or by solving the linear response eigenvalue problem in a time-dependent approach with DFTB methods.
II.4.2 I/O

### v3

@@ Abstract
The software for chemical interaction networks (SCINE) project aims at pushing the frontier of quantum chemical calculations on molecular structures to a new level. While calculations on individual structures as well as on simple relations between them have become routine in chemistry, new developments have pushed the frontier in the field to high-throughput calculations. Chemical relations may be created by a search for specific molecular properties in a molecular design attempt or they can be defined by a set of elementary reaction steps that form a chemical reaction network. The software modules of SCINE have been designed to facilitate such studies. The features of the modules are (i) general applicability of the applied methodologies ranging from electronic structure (no restriction to specific elements of the periodic table) to microkinetic modeling (with little restrictions on molecularity), full modularity so that SCINE modules can also be applied as stand-alone programs or be exchanged for external software packages that fulfill a similar purpose (to increase options for computational campaigns and to provide alternatives in case of tasks that are hard or impossible to accomplish with certain programs), (ii) high stability and autonomous operations so that control and steering by an operator is as easy as possible, and (iii) easy embedding into complex heterogeneous environments for molecular structures taken individually or in the context of a reaction network. A graphical user interface unites all modules and ensures interoperability. All components of the software have been made available open source and free of charge.
@@ I Introduction
The software modules created within the SCINE framework revolve around a common set of concepts and ideas shown in Fig. 1. These ideas have been developed in our research group since 2008. The first concept was interactive and haptic quantum chemistryMarti and Reiher (2009); Haag and Reiher (2013); Haag et al. (2014), which allows a human to manipulate a chemical system and experience its response in real time that can be exploited in the exploration of chemical reaction spaceHaag, Marti, and Reiher (2011). The ultrafast calculations that underlie interactive quantum mechanical studies led us to introduce the concept of molecular propensityVaucher and Reiher (2016) for assessing the likelihood of a reactive system to change a currently explored Born–Oppenheimer surface (e.g., through reduction, oxidation, spin change, and protonation) in order to screen for unexpected reactive events that allow for new discoveries during mechanistic explorations. The general structure of how we explore chemical reaction space has already been sketched in Ref. 13 (see Fig. 1 in that paper). Uncertainty quantificationMortensen et al. (2005); Simm, Proppe, and Reiher (2017); Pernot (2022); Weymuth and Reiher (2024) is an important concept, which we introduced in our quantum chemical reaction exploration framework in order to increase its predictive powerSimm and Reiher (2016, 2018); Reiher (2022). In order to make automated and interactive quantum chemical studies as reliable as possible, we have been working on stable algorithms to facilitate autonomous raw data productionVaucher, Haag, and Reiher (2016); Mühlbach, Vaucher, and Reiher (2016); Vaucher and Reiher (2018), since human intervention in case of troubled or failed calculation would not be possible in such settings. To increase the efficieny of unsupervised automated reaction network explorations, we introduced first-principles heuristics to screen for reactive events in a way that is agnostic with respect to element types as it is based on the interpretation of the electronic wave function of the reactantsBergeler et al. (2015); Grimmel and Reiher (2019, 2021). When brought together, these at first sight somewhat disparate concepts have provided the foundation for the fully automated exploration of complex chemical reaction networks that have now become possible within our SCINE framework. In 2017, we presented the first version of our Chemoton software implementing such explorationsSimm and Reiher (2017), which was the successor of our initial implementation of automated reaction exploration algorithms utilizing first-principles heuristicsBergeler et al. (2015). Starting in 2018, we have begun a major rewrite of the entire software base, implementing a flexible and modular architecture. This resulted in the SCINE framework as it is available today and presented in this work. This flexibility allows one to grow the SCINE framework and incorporate new ideas. The latest additions are the ability to harness human intuition for the efficient navigation in chemical space with a steering wheelSteiner and Reiher (2024), microkinetical modeling to efficiently drive explorations forwardBensberg and Reiher (2023a), and the incorporation of the new concept of lifelong machine-learning modelsEckhoff and Reiher (2023).
Figure 1: Central scientific concepts implemented by the SCINE software modules with corresponding references (in the black boxes).
@@
Within SCINE, we address these challenges by a strictly modular architecture, a loose coupling between the individual modules, and consistent interfaces between them. Individual modules which are only loosely coupled in terms of their source code (i.e., the flow of data between them is clearly defined and limited to only a few instances) are easier to maintain than a complex monolithic architecture. At the same time, this makes developing the software easier and less error-prone, as so-called side effects (i.e., unexpected and unwanted effects upon changing the source code) are largely minimized. Finally, this approach naturally leads to individual modules that are fully functional programs in their own right with useful features. For less complex tasks, it is therefore sufficient to install only a single or a few required modules without having to deal with a particularly complicated software stack. (We note, however, that also the entire SCINE toolchain is straightforward to install by relying on the so-called bootstrap procedure offered by the module Puffin; see below for more details.)
@@ II Overview of SCINE Modules
Figure 3: SCINE modules and their communication. Data flows from and to Heron, Utilities, and Core are not shown for clarity.
@@
The calculations are then carried out by another module called PuffinBensberg et al. (2023c). Puffin retrieves information on new calculations from the database, carries them out, and then writes the results back to the database. For carrying out a calculation, Puffin typically invokes other SCINE modules. For example, it can call SwooseBrunken, Csizi, and Reiher (2021); Brunken and Reiher (2020, 2021) for QM/MM calculations or KiNetXBensberg et al. (2023d); Proppe and Reiher (2019); Bensberg and Reiher (2023b) for microkinetic modeling. For most calculations, however, Puffin relies on ReaDuctBensberg et al. (2023e). The ReaDuct module carries out a wide range of structure-manipulating tasks such as structure minimizations, transition state searches, intrinsic reaction coordinate calculations, B-spline interpolation and optimization of reaction pathsVaucher and Reiher (2018), and Newton trajectory calculationsUnsleber, Grimmel, and Reiher (2022).
@@
Large-scale explorations typically require millions of quantum chemical calculations. Therefore, fast methods with sufficient accuracy are of vital importance. Machine learning potentials can offer both high accuracy and high speed when trained on sufficiently accurate data. However, such potentials require extensive and costly training and often have difficulties learning additional data without “forgetting” previous knowledge. Moreover, most structural descriptors cannot efficiently represent many different atom types. We recently addressed these problems by introducing the new concept of lifelong machine learning potentialsEckhoff and Reiher (2023) implemented in the SCINE module Parrot. Other well-known machine learning potentials such as ANISmith, Isayev, and Roitberg (2017); Devereux et al. (2020), M3GNetChen and Ong (2022), and MACEBatatia et al. (2022) are also available in Parrot.
Another feature which is critical for large-scale quantum chemical calculations is parallelization. Performance-critical parts of the SCINE framework are parallelized with OpenMP. Furthermore, the various parallelization features of the external quantum chemistry packages mentioned above are also supported: When invoking ReaDuct, the user can specify the number of cores which should be utilized. This setting is then forwarded to the underlying software. For example, in the case of ORCA, the number of cores will be written to the corresponding ORCA input file, while in the case of Turbomole, the environment variable PARNODES will be set.
@@
Since Chemoton is based on the first principles of quantum mechanics, it is applicable to any chemical system. In practice, this might require specific quantum chemical methods. For example, for very large systems such as proteins, even fast semi-empirical methods like the ones provided by Sparrow are typically too time-intensive for large-scale explorations. In this case, one can resort to molecular mechanics calculations. To enable such type of modeling out of the box, the SCINE software stack includes a module called Swoose. Swoose can provide energies and forces from molecular mechanics for any chemical system and run molecular mechanics simulations. If no parameters are available for the system, it can automatically parametrize a system-focused force field from quantum chemical reference calculationsBrunken and Reiher (2020). Large systems for which accurate reference calculations would be too costly are automatically fragmented. Moreover, Swoose is also able to carry out QM/MM calculations, in which a key region within a system is automatically defined and then described by an electronic structure model, while the rest of the system is described by a standard force field (supported by machine learning corrections)Brunken and Reiher (2021).
@@
Open-shell systems often require a multi-configurational wavefunction ansatz to be described qualitatively correctly. Multi-configurational methods such as the density matrix renormalization group (DMRG)White (1992, 1993); Chan et al. (2008); Chan and Zgid (2009); Marti and Reiher (2010); Schollwöck (2011); Chan and Sharma (2011); Wouters and Van Neck (2013); Kurashige (2014); Olivares-Amaya et al. (2015); Szalay et al. (2015); Yanai et al. (2015); Baiardi and Reiher (2020) typically require the definition of a so-called active orbital space on input. However, this active space is often not trivial to determine. Our autoCAS algorithmStein and Reiher (2016); Stein, von Burg, and Reiher (2016); Stein and Reiher (2017), implemented in a module of the same nameStein and Reiher (2019); Bensberg et al. (2023f), provides an automated way to reliably determine the active space of any molecule. Moreover, once the active space is determined, autoCAS can also invoke software to carry out multi-configurational calculations. Besides QCMaquisKeller et al. (2015), which is our software for DMRG calculations and driven by autoCAS via OpenMolcasLi Manni et al. (2023), autoCAS is also interfaced to ChronusQWilliams-Young et al. (2020), PySCFSun et al. (2020), and SerenityUnsleber et al. (2018). We note that QCMaquis can do much more than just calculate the ground state electronic energy of a molecule. It can also optimize excited statesBaiardi et al. (2019); Baiardi, Kelemen, and Reiher (2022), solve the vibrational Schrödinger equation for anharmonic potential energy surfacesBaiardi et al. (2017); Glaser, Baiardi, and Reiher (2023), and carry out quantum dynamics simulationsBaiardi and Reiher (2019); Baiardi (2021). Furthermore, it features embedding approachesDresselhaus et al. (2015); Hedegård and Reiher (2016), a non-orthogonal state interaction methodKnecht et al. (2016), relativistic calculationsBattaglia, Keller, and Knecht (2018), multi-component DMRGMuolo et al. (2020); Feldmann et al. (2022), and N-electron valence state perturbation theory calculationsFreitag et al. (2017).
@@
A central module is HeronBensberg et al. (2022), our graphical user interface. It aims to make the entire SCINE software ecosystem easily accessible. Heron provides the ability to browse the contents of a database, visualize and analyze existing reaction networks, and operate Chemoton (for example, with the so-called Steering WheelSteiner and Reiher (2024)). Moreover, it is a graphical user interface for Swoose, ReaDuct, autoCAS, and Sparrow. Together with the latter, it provides the ability to interactively carry out quantum chemical calculations, in which the response of a molecule to structural manipulations can be experienced in real time. This can even be coupled with haptic devices, allowing a person to literally feel the forces acting on atoms in a moleculeMarti and Reiher (2009); Haag and Reiher (2013); Haag et al. (2014); Weymuth and Reiher (2021).
@@ II.1.1 Field of application
MolassemblerBensberg et al. (2023b); Sobez and Reiher (2020) provides characterization, manipulation, and conformer generation for organic and inorganic molecules. In principle, it can handle any organic or inorganic molecule, including haptic and multidentate ligands and almost arbitrary coordination spheres. Its general approach to the characterization of molecules and its broad applicability makes it distinct compared to similar programs such as RDKitRDK (2023) or openBabelO’Boyle et al. (2011), which focuses on organic molecules, or molSimplifyIoannidis, Gani, and Kulik (2016); Edholm et al. (2023), AARONGuan et al. (2018) and DENOPTIMFoscato, Venkatraman, and Jensen (2019), which provide utilities to design new molecules, or CRESTPracht, Bohle, and Grimme (2020) for conformer generation.
@@
Molassembler characterizes a molecule’s connectivity by a graph and the local geometry of each atom by polyhedral shapes. In organic chemistry, atoms typically have only up to four bonding partners, leading at most to a tetrahedral local environment. By contrast, transition metal atoms can have significantly more neighbors and diverse coordination geometries. Therefore, Molassembler supports polyhedra for the local geometry up to icosahedra and cuboctahedra and can be applied to the full spectrum of organic and inorganic chemistry. Furthermore, it supports multidentate and haptic ligands. Molassembler differentiates stereoisomers through a ranking algorithm based on generalizing the ranking rules for organic molecules provided by the International Union of Pure and Applied ChemistryFavre and Powell (2013). The abstract molecule representation provided by Molassembler can be manipulated to change the connectivity in the molecule and replace or edit ligands to generate new molecules while retaining any chiral information in the remaining molecule wherever possible. Furthermore, Molassembler can generate conformers (either by by enumerating the rotamers of the molecule or by relying on Distance GeometryJeffrey M. Blaney (2007) for a stochastic conformer generation), generating initial Cartesian coordinates for the conformer guesses before refining the coordinates.
II.1.2 Scientific background
In the context of SCINE, one task delegated to Molassembler is the interpretation of the results of quantum chemical calculations into molecular graphs. Because the input structures to these calculations may be of unknown stability or of purposeful instability such as in the case of reactivity explorations, no information regarding the bond connectivity or stereochemistry of the input structure may be transferred onto the output structure of the calculation. Due to these limitations, the interpretation of a structure proceeds with only its atom-labeled spatial positions, and possibly a quantum chemically calculated bond order matrix, as inputs.
Prior to the inference of stereochemical information, possibly available bond order information is combined with internuclear distances into connectivity. The resulting graph can be split into its connected components, each of which represent either a molecule or a singular atom.
Molassembler represents atom-centered stereochemistry as a discrete local polyhedral shape combined with a surjection from shape vertices to the chemically ranked ligand binding sites. There are a number of subtleties to this representation, of which it is sufficient here to mention only one: The discretization of varied local coordination geometries into idealized polyhedral shapes necessarily limits the capacity to represent structures other than ground-state minima, since transition states or excited states can exhibit significantly irregular local shapes.
The classification of an arbitrary local coordination geometry or bonding pattern into discrete shapes is settled on a protocol where the distance between input and shape is quantified by a continuous shape measurePinsky and Avnir (1998). The time-determining step of the algorithm involves finding a point to vertex matching that minimizes a quaternion fit. The implementation in Molassembler reduces the factorial complexity of this step by exploiting rotational symmetries of the target shape and a heuristic considering a rotational fit sufficiently converged with five vertex matchings. This allows for calculations with shapes up to twelve vertices in size on the order of seconds. Additionally, the origin of the spherical shapes is pre-matched to the atomic nucleus at the center of the local geometry, adding more information into the quaternion fit. A complication stemming from the incomparability of continuous shape measures with respect to different target shapes has been solved by further calculating the probability that each measure is the result of a random input geometry. The shape whose continuous shape measure is least likely to have been calculated from a random input geometry is selected. We refer to the original publication in Ref. 34 for further details on the algorithms implemented.
From the point to vertex matching gained from selecting a local shape and chemically ranking ligand and substituent binding sites, local stereoinformation can be constructed.
II.1.3 API
@@ II.4.1 Field of application
SparrowBosia et al. (2023a); Bosia, Weymuth, and Reiher (2022); Bosia et al. (2023b) specializes in ultra-fast quantum chemical calculations with semi-empirical methods, which are essential in applications such as interactive quantum chemistry or large-scale reaction network exploration. Calculations with semi-empirical methods based on the neglect of diatomic differential overlap approximation — NDDO methods — as well as with methods based on the density functional tight binding approach — DFTB methods — are available (see below). Ground state calculations yield electronic energies, Hessian matrices, nuclear gradients, bond orders, and thermodynamic properties such as Gibbs free energies. Electronically excited states can also be calculated — in a configuration interaction approach with NDDO methods or by solving the linear response eigenvalue problem in a time-dependent approach with DFTB methods.
Through the interactive user interface HeronBensberg et al. (2022), an operator is able to interact with a molecular system of interest in real time. This is made possible by real-time quantum chemical calculationsHaag and Reiher (2013) implemented as ultrafast semi-empirical approaches in Sparrow. Specifically, the molecular structure(s) relax toward an energy-minimum structure while the operator may manipulate the structure simultaneouslyHaag and Reiher (2014); Vaucher, Haag, and Reiher (2016); Vaucher and Reiher (2017) (e.g., in search of viable reaction pathsHaag, Marti, and Reiher (2011)). For a smooth operator experience of these manipulations, a high-frequency continuous feedback of electronic energies and gradients of at least 25 Hz is requiredMark et al. (1996); Ruspini, Kolarov, and Khatib (1997).
We emphasize that Heron does not only offer an interactive experience with the computer mouse, but further allows the operator to feel the quantum chemical forces on atomic nuclei in the context of haptic quantum chemistryMarti and Reiher (2009). That is, the gradient of the energy at the position of a selected atom can be experienced by an operator’s tactile sense as attraction or resistance along the direction of manipulation through a force-feedback haptic deviceMarti and Reiher (2009).
II.4.2 Scientific background
Sparrow implements semi-empirical methods of the neglect of the diatomic differential overlap (NDDO) and density functional tight binding (DFTB) families. These allow for ultrafast, albeit approximate quantum chemical calculations.
Within the NDDO approximation, the Hartree–Fock formalism is modified by neglecting the differential overlap between atomic orbitals $\chi_{\mu^{I}}$ and $\chi_{\nu^{J}}$ on different atomic centers, i.e., $\delta_{IJ}\chi_{\mu^{I}}\chi_{\nu^{J}}\approx\phi_{\mu}\phi_{\nu}$, where $\chi_{\mu^{I}}$ denotes the basis function $\mu$ of an atomic orbital expansion in a non-orthogonal basis centered on the atom $I$, and $\phi_{\mu}$ is the corresponding basis function in an orthogonal basis, such that
$\phi_{\nu}=\sum_{\mu}^{M}\left(S^{-\frac{1}{2}}\right)_{\mu\nu}\chi_{\mu}\;.$ (1)
As a consequence, the two-electron repulsion integrals in an orthogonalized atomic orbital basis can be approximated in terms of two-electron integrals in a non-orthogonal atomic orbital basis as
$\langle\phi_{\mu}\phi_{\nu}|\phi_{\lambda}\phi_{\sigma}\rangle\approx\delta_{IJ}\delta_{KL}\langle\chi_{\mu^{I}}\chi_{\nu^{J}}|\chi_{\lambda^{K}}\chi_{\sigma^{L}}\rangle\;.$ (2)
In all NDDO methods, all three- and four-center two-electron repulsion integrals are neglected. The advantage is twofold: First, their computational expensive calculation is avoided, and second, the NDDO Fock matrix ${}^{\textrm{NDDO}}\bm{F}$ assembled from these integrals is already expressed in an approximately orthogonal basis. NDDO methods solve the eigenvalue problem
${}^{\textrm{NDDO}}\bm{F}{}^{\textrm{NDDO}}\bm{C}={}^{\textrm{NDDO}}\bm{\varepsilon}{}^{\textrm{NDDO}}\bm{C}\;,$ (3)
where the ordinary Fock, molecular orbital, and diagonal orbital energy matrices are substituted with their NDDO counterparts. The overlap matrix is set to be equal to the identity matrix by virtue of the orthogonalized basis in which the NDDO calculation is carried out. The specifics of the calculation of the Fock matrix are heavily method-dependent and can be found in Ref. 54.
By contrast, DFTB methods are constructed by expanding the energy functional with respect to the electronic density to various degrees of a Taylor series expansion of $\rho$ around a a reference density $\rho_{0}$,
$\begin{split}{}^{\textrm{DFTB}}E[\rho_{0}+\delta\rho]&=E^{(0)}[\rho_{0}]+E^{(1)}[\rho_{0}+\delta\rho]\\ &+\frac{1}{2}E^{(2)}[\rho_{0}+\left(\delta\rho\right)^{2}]+\frac{1}{6}E^{(3)}[\rho_{0}+\left(\delta\rho\right)^{3}]\\ &+...\quad.\end{split}$ (4)
The DFTB0, DFTB2, and DFTB3 models include terms up to the first, second, and third order, respectively. DFTB0 calculations do not require iterative solutions until self-consistency has been reached; they include the zeroth and first order of the energy functional expansion. $E^{(1)}[\rho_{0}+\delta\rho]$ is obtained by solving the eigenvalue problem
[…]
$E^{(1)}[\rho_{0}+\delta\rho]=\sum_{i}^{M}n_{i}\varepsilon_{i}\quad,$ (6)
with $M$ being the number of electrons in the system. $E^{(0)}[\rho_{0}]$ is approximated as a sum of diatomic repulsion terms between atoms $I$ and $J$, $V_{\textrm{rep},IJ}$, modeled to correct the sum of the occupied molecular orbital energies with respect to DFT reference calculations. The resulting DFTB0 energy reads
${}^{DFTB0}E[\rho_{0}+\delta\rho]=\sum_{i}^{M}n_{i}\varepsilon_{i}+\frac{1}{2}\sum_{I}\sum_{J}V_{\textrm{rep},IJ}\quad.$ (7)
DFTB2 includes the effects of a self-consistent redistribution of Mulliken charges to the energy functional and the elements of the Hamiltonian operator. The corresponding energy expression is given by
${}^{DFTB2}E^{(2)}$ $=\sum_{I}\sum_{J}\Delta q_{I}\Delta q_{J}\gamma_{IJ}$ (8)
${}^{DFTB2}\mathcal{H}_{\mu^{I}\nu^{J}}$ $=\mathcal{H}^{(0)}_{\mu^{I}\nu^{J}}+\frac{S_{\mu^{I}\nu^{J}}}{2}\sum_{K}^{N}\Delta q_{K}\left(\gamma_{IK}+\gamma_{JK}\right)\quad,$ (9)
where the notation $\mu^{I}$ indicates the atomic orbital $\mu$ located on the atomic center $I$, $\Delta q_{I}$ is the Mulliken charge on atom I, $\gamma_{IJ}=\frac{1}{R_{IJ}}-S$, and $S$ is a model parameter. DFTB3 introduces another term into the energy density functional expansion. The energy correction and the Hamiltonian coefficients are then given by
$\begin{split}{}^{DFTB3}E^{(3)}&=\sum_{I}\sum_{J}\Delta q^{2}_{I}\Delta q_{J}\Gamma_{IJ}\;,\\ {}^{DFTB3}\mathcal{H}_{\mu^{I}\nu^{J}}&={}^{DFTB2}\mathcal{H}_{\mu^{I}\nu^{J}}+\frac{S_{\mu^{I}\nu^{J}}}{3}\sum_{K}^{N}\Delta q_{K}\Gamma_{IJK}\;,\\ \Gamma_{IJK}&=\left(\Gamma_{IK}\Delta q_{I}+\Gamma_{JK}\Delta q_{J}+\frac{\Delta q_{K}}{2}(\Gamma_{IK}+\Gamma_{JK})\right)\;,\end{split}$ (10)
where the function $\Gamma_{IJ}$ represents the derivative of $\gamma_{IJ}$ with respect to the atomic charge, and DFTB3 includes in $\gamma_{IJ}$ another correction to improve on the performance for certain systems.
The semi-empirical methods available in Sparrow provide fast and reliable electronic energies and gradients. However, the calculation cost still scales with the number of atoms. Therefore, to guarantee the required high frequency for seamless visual and haptic feedback, the potential energy surface is approximated for atomic positions between finalized quantum chemical calculation with a Taylor series expansion around the atomic coordinates of the last finished electronic structure calculation; the resulting harmonic potential is called mediator or surrogate potentialVaucher, Haag, and Reiher (2016). The mediator potential resembles the local Born–Oppenheimer surface and prohibits escape to regions of configuration space for which no electronic structure information has become available yet.
This real-time interaction is of interest to both researchers who would like to explore their reactive system in a more intuitive and interactive wayCsizi, Steiner, and Reiher (2023) as well as for learners who are in search for an immersive experience of computational chemistryMüller, Reiher, and Kapur (2024).
II.4.3 I/O